# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ 4.1.4. Replanning
After reverting back to the original step description, Nóírín is curious about authors who have published on “interactive feedback mechanisms in AI” rather than seeing papers directly. She edits the step description to reflect this and saves it. At this point, the plan’s trajectory has changed—the next step, which has the agent suggest common themes in papers, is now irrelevant because none of the previous steps retrieve papers. The system automatically detects this and replans subsequent steps accordingly by removing those steps and “autocompleting” the plan with new ones. Just like with full plans, Cocoa will present a selection of new steps for the user to choose from. Nóírín selects an option where the agent searches for some papers by the authors it found earlier in the plan before summarizing some common themes across collected papers.
Figure 6. If major changes are made to a step that changes the rest of the plan’s trajectory, Cocoa detects this and will trigger replanning. Replanning replaces subsequent steps with ones that the agent suggested for “autocompleting” the plan.
@@ 4.2.1. Continuous and stepwise execution
Nóírín can trigger continuous co-execution with the [Run all] button (or an [Update and run] button if the plan was edited) at the top of the plan. In continuous co-execution, the agent will automatically continue onto the next step as soon as it (or the user) has completed the previous one, until it reaches a user step or the end of the plan.
By contrast, stepwise co-execution allows Nóírín to run one plan step at a time, taking inspiration from computational notebooks. When Nóírín hovers her mouse over the bullet point of the first plan step, the bullet point turns into a play button that, when clicked, will run only that step. Upon the step’s completion, the agent will not execute the next step until Nóírín manually clicks that step’s play button. Unlike some computational notebooks, stepwise execution does not permit out-of-order execution of steps [^3] We also note that out-of-order execution in computational notebooks is a major user pain point identified in prior academic work (Chattopadhyay et al., 2020; Lau et al., 2020) and industry reports (Jupyter, 2015; Grus, 2018). because some steps may rely on the outputs of previous ones. The interface reinforces this by only allowing a step’s bullet point to turn into a play button on hover if the previous step has been completed, or if it is the first step. Nóírín can switch from stepwise to continuous co-execution at any point by clicking the [Run remaining] button at the top of plan. Conversely, Nóírín can switch from continuous to stepwise co-execution by clicking on the [Pause after this step] button during execution and manually run subsequent steps.
If Nóírín wants to rerun any completed step, she can easily do so by hovering over the step’s checkmark indicator and clicking on the play button that appears. Knowing this, Nóírín initiates continuous co-execution of her edited plan by clicking on the [Update and run] button.
@@ 4.2.2. Completing user steps
The agent completes the first step of the plan and arrives at a user step. Here, the agent requires user input to continue. Nóírín can easily see this because of the orange highlighting on the step for which her input is needed. She clicks into the text that appeared under the step prompting her to provide her guidance and opens an interactive sidebar. The sidebar offers her a built-in paper search functionality connected to the Semantic Scholar academic database and she uses this to add some relevant papers she previously encountered. She clicks Save, adding the papers and their metadata to a plan-specific context pool that the agent references when completing future steps. The agent then proceeds to complete the next step with this updated context to expand upon Nóírín’s selected papers.
@@ 4.2.3. Editing outputs of agent steps
The agent has now completed the third step, where it conducted a paper search guided by Nóírín’s seed papers from the previous step. Nóírín clicks into a text preview of the output that appeared under the step, opening a sidebar similar to the one she used in the previous step to add papers, this time populated with interactive paper cards. This presents opportunities for her to further guide the agent with her expertise by removing papers she considers irrelevant and adding papers the agent might have missed. She sees a couple of low-quality results that have incoherent titles and removes them by clicking X on the card. She also uses the built-in topic search in the sidebar to search for and add topics to the list. After a couple different searches, she adds one more paper she recalls from a past discussion with a collaborator and saves this curated list. The papers in this list, along with its metadata, are stored in the plan’s shared context pool.
@@
The interactive UI in the sidebar dynamically adapts according to the step’s output type, which fall into one of categories in Cocoa: papers, authors, topics, entities, and text. Papers, authors, and topics are drawn from the Semantic Scholar and are displayed as interactive cards that Nóírín can add or delete. The sidebar has built-in Semantic Scholar search for papers, authors, and topics, the results from which she can add to the agent’s output. Entities are lists of items in natural language (search queries, research questions, etc.) and are displayed as editable pills. Text is natural language text displayed in an editable text box.
@@
The system tracks how many times Nóírín edits agent outputs as an indicator of output reliability. If she edits an output more than twice on a particular agent step, the agent will pause and seek her confirmation on that step before continuing. Moreover, if the agent fails to return any output, the system will alert Nóírín and ask her to complete the step instead. These features loop in human guidance to bolster the quality of outputs in the face of weak agent performance or agent failures.
@@
Figure 8. Users can click the output preview text beneath the plan step to access the interactive sidebar (1). By editing agent outputs in the interactive sidebar (2), users draw from their expertise, such as papers the agent might have missed (3), to guide the agent. Once edited, an indicator will appear in the output’s text preview below the step (4).
@@
After the final plan step has been completed, the agent adds a plan output panel (Fig. 9) containing a modified version of the last step’s output into the document just below the plan. The modification involves rewriting the output, given the context of the plan and its outputs, to more directly connect to the original user request. This brings the results of running the plan into the document so Nóírín can easily reference it in the context of other content. The panel is collapsed by default to save space and reduce clutter, but she can expand it to copy and paste parts of the output she finds useful and add them to her document. If she does not want the panel in her document at all, she can simply delete it; she can still access the outputs of all steps in the sidebar.
@@
Figure 9. Once the execution of the entire plan is complete, the agent inserts a plan output panel in the document with a modified version of the last step’s output. This allows the user to view the plan’s results within the context of their document. The panel can be expanded or collapsed, and remains in the document even if the plan itself is collapsed for easy access.
4.3. Iterative Co-Planning and Co-Execution
@@
In addition to supporting co-planning and co-execution independently, Cocoa is designed to support synergistic blending of the two. As Nóírín co-executes the plan with the agent, she encounters an author that the agent identified who publishes highly relevant work and wishes to further explore the author’s papers. She edits the following step’s plan description to satisfy this. She reruns that step and receives a new batch of papers, and the agent automatically reruns the rest of the plan with the updated output. She is also not satisfied with the agent’s interpretation of common themes and believe that she can surface deeper insights by reading the papers herself, so she toggles that step to be a user step. After having some time to read the papers, she returns to her document and jots down her insights. In doing so, has already started reasoning about new feedback elicitation mechanisms for agents. She now finds the next step repetitive, so she deletes it. She reruns the last step on her updated notes to see the agent’s constructive critiques. Overall, her use of Cocoa is highly iterative, and she smoothly transitions back-and-forth between co-planning and co-execution.
So far, Nóírín has just been interacting with one plan. While waiting for the agent to complete a series of steps, she uses the [Collapse] button to hide the steps and only reveal essential information about the plan’s status. She identifies a couple more areas of her document that merit further exploration and invokes the agent on them. Soon, she is co-planning and co-executing with multiple agents on different parts of her document in parallel. This is not cognitively burdensome for her as she only has to attend to one plan at a time as the others are all collapsed and running in the background (See Fig. 10).
@@ 4.4.1. Technical stack
Cocoa is implemented as a web application with a Next.js and TypeScript frontend communicating with a Flask backend. The frontend uses the Tiptap [^4] Tiptap is an open-source headless wrapper for ProseMirror, a toolkit for building web-based rich text editors. framework for the main document editor. Each document supports synchronous collaboration via Hocuspocus [^5] Hocuspocus is a Y.js WebSocket backend for conflict-free real-time collaborative web apps. and all changes are auto-saved to Tiptap Cloud, which also serves as storage for all participant documents. Interactive components within the document editor are implemented as custom Tiptap extensions in TypeScript.
@@ 4.4.2. Underlying LLM Agent
The Flask backend orchestrates LLM activity with calls to GPT-4o (scaffolded by LangChain [^6] https://www.langchain.com/). To optimize for speed, simple actions such as plan generation or text summarization, the system makes a direct call to GPT-4o. For plan steps that required access to the scientific literature, we created a custom tool-calling LLM agent powered by GPT-4o that had two tools that allowed it to access the public Semantic Scholar API [^7] https://www.semanticscholar.org/product/api, which covers paper/author/topic search, and the Ask this Paper [^8] https://www.semanticscholar.org/product feature for paper summarization and question-answering. While GPT-4o provides APIs for managing multi-turn conversation context, the system manages its own context and always calls GPT-4o single-turn completion API. All prompts used are provided in Appendix F.
@@ 4.4.3. Guidance for plan generation
We use findings from our formative study to guide the agent’s generation of initial plans. We had asked participants to write down brief plans as part of the probe activity (see Section C.2) and affinity mapped participants’ plan steps into themes, after which we created a “step template” for the theme. For example, many participants wrote a paper search step, for which we created the following template: Search for papers that discuss [query] and sort by [criteria]. Based on participants’ preferences of which steps they prefer assigning to an AI vs. keeping for themselves (RQ3 from our formative study), we organized these steps into agent and user steps. We then wrote examples of how these steps are composed into plans for tackling particular questions, once again drawing from plans participants wrote in the probe activity. These examples were provided for in-context learning in Cocoa’s system prompt for plan generation (see Appendix F).
@@
5. User Study
We conducted a within-subjects user study with 16 researchers to evaluate the effectiveness of Cocoa—and by extension, our design pattern of interactive plans—against a strong chat baseline. Specifically, our study aimed to address the following research questions:
  • RQ1:
@@ 5. User Study
  • RQ2:
    When did researchers prefer interacting with an AI agent through interactive planning over our chat baseline, and vice versa?
  • RQ3:
@@ 5.1. Participants
We recruited 16 Ph.D. and postdoctoral researchers (14 Ph.D.s, 2 postdocs; 10 female, 6 male) in computer science (CS) or CS-adjacent areas via university mailing lists, word of mouth, social media recruitment messages (on Twitter/X, Mastodon, Bluesky), and personal connections. One participant also participated in our formative study. 15 researchers were based in the United States, and 1 was based in Canada. Participants’ research areas ranged from large language models, to ubiquitous computing, to visualization; broadly, they spanned HCI ($n=10$), ML ($n=3$), and NLP ($n=3$). We sent out a recruitment form to collect basic demographic details, frequency of AI use in research, and a copy of a project document from an ongoing or completed project document to use during the study. We recruited on a first-come, first-served basis as we conducted our studies and closed recruitment when we approached data saturation. In our participant pool, many users were occasional (6) or frequent [^9] We define “occasional” and “frequent” the same way as we do in our formative study (Table 1). (6) users of AI tools in the research process. We also recruited 4 additional participants (1 female, 2 male, 1 non-binary) from our institution for pilot studies to test out Cocoa, catch usability issues, and help us refine our procedure. We learned from our pilot studies that the length of generated plans should be around 3 steps (rather than the system’s default of 5–6 steps) to fit within the study’s time constraints. Further details of our study participants can be found in Table 2 of Appendix D.
@@ 5.2. Baseline System
The baseline system (Figure 11) was a chat interface powered by the same LLM agent used in Cocoa: it used GPT-4o with access to the same tools for completing literature-related tasks using a combination of Semantic Scholar and standard LLM capabilities (summarization, accessing knowledge stored in weights, etc.). While we could integrate interactive plans within chat and vice versa (more on this in Section 7.1), we opted not to for a cleaner comparison of the two design patterns. The design of the chat interface closely resembled that of popular LLM chatbots such as ChatGPT, Claude, and Gemini. This chat interface was used alongside the same document editor as the one Cocoa uses, except we removed the option to invoke the agent, thereby eliminating all interactions related to co-planning and co-execution. During the study, participants positioned the chat interface beside the baseline document editor for easy access.
@@
Figure 11. Participants’ typical setup for our baseline condition. On the left, the same document editor from Cocoa but without an option to invoke the agent for co-planning and co-execution. On the right, the agent situated within a chat interface.
5.3. Research Task
@@
Since the capability of our underlying agent focused on helping users explore and understand scientific literature, we designed our study so that our participants tackled a literature-augmented task in each of Cocoa and the baseline. By literature-augmented, we mean tasks that are to be completed by referencing or drawing from academic literature; in the study, participants worked on not only literature review tasks, but also tasks where researchers need to make literature-informed decisions such as study design and ideation. This stands in contrast with tasks focused on writing mechanics (e.g., rewording a sentence) or tasks that do not involve literature (e.g., booking travel) that past work has covered (e.g., (Yao et al., 2022a; Yang et al., 2024; Kim et al., 2023; Zhang et al., 2023; Ma et al., 2023; Lee et al., 2024a)).
To make the study more realistic and grounded in real-world research projects, tasks for this study were open questions or unfinished items from participants’ existing project documents. To control for the length of the study, the first author reviewed the participants’ documents prior to the study to identify and highlight two candidate tasks. To make sure the two tasks are valid and comparable, early on in the study, the first author asked participants 1) whether the tasks have been decisively solved by the participant and whether they would still like to explore them in the study, and 2) whether the two tasks are similar in scope, such that one task does not take significantly more resources and effort to explore than the other. If these conditions were not met, the participant was prompted to rewrite one or both tasks until they were. The two tasks were then randomly assigned to Cocoa and the baseline.
5.4. Procedure
The first author (study facilitator) conducted 1:1 studies with the 16 participants over Google Meet between October and December 2024. Each study was 90 minutes in length. The study started off with introductions and a brief overview, followed by the two tasks, which were counterbalanced across participants.
In the Cocoa condition, participants first watched a 6-minute video demo of the system, and were then directed to a document where they would get used to using Cocoa’s features on the following sample question: “How can we use AI to better society?” This practice period lasted around 10 minutes. Participants were then asked to spend 25 minutes to make as much progress as possible tackling the task assigned to this condition with Cocoa, thinking aloud as they did so. They were also permitted to also use other tools (academic search engines, other AI tools, social media, etc.) alongside Cocoa, although very few did. After the 25 minute session, participants wrote a brief set of takeaways and next steps from their exploration, and filled out a short evaluation form with 5-point Likert scale questions about their experience (see Appendix E for this form). After submitting the form, they were encouraged to further try out Cocoa on other parts of their project document for another 10–15 minutes.
@@ 5.4. Procedure
The study concluded with a semi-structured interview that lasted around 15 minutes. Participants were asked to reflect on their experiences across the two systems and discuss the pros and cons of both. They were also asked about particular decisions the study facilitator observed when using each system. After this study concluded, participants received a \$75 USD honorarium for their participation. All studies were recorded and transcribed by Google Meet. This study was reviewed and approved by our organization’s internal IRB.

### v3

@@ 4.1.4. Replanning
At this point, Nóírín’s edits have changed the plan’s trajectory—the next step, which has the agent suggest common themes in papers, is now irrelevant because none of the previous steps retrieve papers. The system automatically detects this and replans subsequent steps accordingly by removing those steps and “autocompleting” the plan with new ones (Figure 6). However, before making these changes, Cocoa notifies Nóírín and asks whether she would like to proceed with replanning; Nóírín accepts. Just like with full plans, Cocoa will present a selection of new steps for Nóírín to choose from. She selects an option to complete the plan.
Figure 6. If major changes are made to a step that changes the rest of the plan’s trajectory, Cocoa detects this and will trigger replanning. Replanning replaces subsequent steps with ones the agent suggested for “autocompleting” the plan.
[…]
The user has edited a step significantly, such that it changes the rest of the plan’s trajectory. Cocoa detects this and replaces subsequent steps with ones the agent suggested for autocompleting the plan.
@@ 4.2.1. Continuous and stepwise execution
Nóírín wishes to run through the entire plan once to see the output. She triggers continuous co-execution (Figure 7A) with the [Run all] button at the top of the plan. In continuous co-execution, the agent will automatically continue onto the next step as soon as it (or the user) has completed the previous one, until it reaches a user step or the end of the plan.
When she iterates on the plan in subsequent runs to refine specific steps and outputs, Nóírín may opt for stepwise co-execution (Figure 7B). This allows Nóírín to run one plan step at a time, like running individual cells in computational notebooks. When Nóírín hovers her mouse over the bullet point of the first plan step, the bullet point turns into a play button that, when clicked, will run only that step. Unlike some computational notebooks, however, Cocoa does not permit out-of-order stepwise execution [^2] We also note that out-of-order execution in computational notebooks is a major user pain point identified in prior academic work (Chattopadhyay et al., 2020; Lau et al., 2020) and industry reports (Jupyter, 2015; Grus, 2018). because some steps may rely on the outputs of previous ones.
Nóírín can flexibly switch from stepwise to continuous co-execution at any point by clicking the [Run remaining] button at the top of the plan. Conversely, Nóírín can switch from continuous to stepwise co-execution by clicking on the [Pause after this step] button during execution and manually running subsequent steps.
@@
[…]
Continuous and stepwise execution are shown side by side. A single button click at the top of the plan triggers continuous execution, in which the agent will automatically move onto the next step as soon as it has completed the previous one. To trigger stepwise execution, users hover over the play button beside each step. The agent will not move on to the next step until the user explicitly clicks the play button for that step.
@@ 4.2.2. Completing user steps
The agent completes the first step of the plan and arrives at a user step that Nóírín assigned herself. Here, the agent requires Nóírín’s input to continue, and Nóírín can see this because the step has been highlighted in orange. When she clicks into that step, an interactive sidebar opens to the right (Figure 8). The sidebar offers her a built-in paper search functionality connected to the Semantic Scholar academic database, and she uses this to add some relevant papers she has in mind. She clicks Save, adding the papers and their metadata to a plan-specific context pool that the agent references when completing future steps. The agent then proceeds to complete the next step with this updated context to perform a more extensive paper search given the seed papers Nóírín provided.
@@ 4.2.3. Editing outputs of agent steps
The agent has now completed the third step, where it conducted a paper search guided by Nóírín’s seed papers from the previous step. Nóírín clicks into that step, opening a sidebar similar to the one she used in the previous step to add papers, this time populated with interactive paper cards. This presents opportunities for her to further guide the agent with her expertise by removing papers she considers irrelevant and adding papers the agent missed. She sees a couple of low-quality results that have incoherent titles and removes them. She then adds one paper she recalls from a past discussion with a collaborator and saves this curated list. The papers in this list, along with their metadata, are stored in the plan’s context pool, accumulated across all completed steps.
@@
The interactive UI in the sidebar dynamically adapts according to the step’s output type, which fall into one of categories in Cocoa: papers, authors, topics, entities, and text. Papers, authors, and topics are drawn from Semantic Scholar and are displayed as interactive cards that Nóírín can add, delete, or search for using the sidebar’s built-in Semantic Scholar search. Entities are lists of items in natural language (search queries, research questions, etc.) and are displayed as editable pills. Text is natural language text displayed in an editable text box.
@@
For any agent step, if the agent fails to return any output, the system will alert Nóírín and ask her to complete the step instead. These features loop in human guidance to bolster the quality of outputs in the face of weak agent performance or agent failures.
@@
Figure 8. Users can click the output preview text beneath the plan step to access the interactive sidebar (1). By editing agent outputs in the interactive sidebar (2), users draw from their expertise to edit outputs, such as adding papers the agent might have missed (3), to guide the agent. Once edited, an indicator will appear in the output’s text preview below the step (4).
[…]
A user clicks the output preview text beneath the plan step to access the interactive sidebar that contains some card UIs displaying paper information. Users can delete card UIs they consider irrelevant. By editing agent outputs in the interactive sidebar, users draw from their expertise to guide the agent. Once edited, an indicator will appear in the output’s text preview below the step.
4.2.4. Plan output panel
@@
After the final plan step has been completed, the agent adds a plan output panel (Figure 9) containing a modified version of the last step’s output into the document just below the plan. The modification involves rewriting the output, given the context of the plan and its outputs, to more directly connect to the original user request. This brings the results of running the plan into the document so Nóírín can easily reference it in the context of other content. The panel is collapsed by default to save space and reduce clutter, but she can expand it to copy and paste parts of the output she finds useful and add them to her document. If she does not want the panel in her document at all, she can simply delete it; she can still access the outputs of all steps in the sidebar.
@@
Figure 9. Once execution of the entire plan is complete, the agent inserts a plan output panel in the document with the last step’s output. This allows the user to view the plan’s results within the context of their document. The panel can be expanded or collapsed, and remains in the document even if the plan itself is collapsed for easy access.
[…]
A light blue panel appears at the bottom of a plan once it’s finished executing to allow the user to view the plan’s results within the context of their document. Content from the panel may be copied by the user and pasted elsewhere in the document. The panel can be expanded or collapsed, and remains in the document even if the plan itself is collapsed for easy access.
4.3. Interleaving Co-Planning and Co-Execution
@@
In addition to supporting co-planning and co-execution independently, Cocoa is designed to support synergistic interleaving of the two. As Nóírín co-executes the plan with the agent, she encounters a relevant author that the agent identified and wishes to further explore the author’s papers. She edits the following step’s plan description to satisfy this. She reruns that step and receives a new batch of papers, and the agent automatically reruns the rest of the plan with the updated output. She is also not satisfied with the agent’s interpretation of common themes and believes that she can surface deeper insights by reading the papers herself, so she toggles that step to be a user step. After reading the papers in detail, she returns to her document and jots down her insights. She now finds the next step repetitive, so she deletes it. Overall, her use of Cocoa is highly iterative, and she smoothly transitions between co-planning and co-execution.
@@
[…]
An illustration of three plans in a document, with two of them collapsed so that only essential information about their operating status is shown without details about the individual steps.
So far, Nóírín has just been interacting with one plan. While waiting for the agent to complete a series of steps, she can collapse the plan to hide the steps and only reveal essential information about the plan’s status. She identifies a couple more areas of her document that merit further exploration and invokes the agent on them. Soon, she is co-planning and co-executing with multiple agents on different parts of her document in parallel. This is not cognitively burdensome for her as she only has to attend to one plan at a time, since the others are all collapsed and running in the background (Figure 10). Finally, the document in Cocoa supports real-time multi-user editing, so Nóírín’s collaborators can also leverage the agent and interact with plans.
@@ 4.4.1. Technical stack
Cocoa is implemented as a web application with a Next.js and TypeScript frontend communicating with a Flask backend. The frontend uses the Tiptap [^3] https://tiptap.dev/. framework for the main document editor. Each document supports synchronous collaboration via Hocuspocus [^4] https://tiptap.dev/docs/hocuspocus/. and all changes are auto-saved to Tiptap Cloud, which also serves as storage for all participant documents. Interactive components within the document editor are implemented as custom Tiptap extensions in TypeScript.
@@ 4.4.2. Underlying LLM Agent
The Flask backend orchestrates LLM activity with calls to GPT-4o (scaffolded by LangChain [^5] https://www.langchain.com/). To optimize for speed, simple actions such as plan generation or text summarization involve a direct call to GPT-4o. For plan steps that required access to the scientific literature, we created a custom tool-calling LLM agent powered by GPT-4o that had two tools that allowed it to access the public Semantic Scholar API [^6] https://www.semanticscholar.org/product/api, which covers paper/author/topic search, and the Ask this Paper [^7] https://www.semanticscholar.org/product feature for paper summarization and question-answering. While GPT-4o provides APIs for managing multi-turn conversation context, the system manages its own context, which includes previous plan steps and their outputs, and calls GPT-4o’s single-turn completion API. All prompts used are provided in Appendix I.
@@ 4.4.3. Guidance for plan generation
We use findings from our formative study to guide the agent’s generation of initial plans. In our formative study, we had asked participants to write down brief plans when interacting with our probe (see Appendix C). Based on participants’ preferences of which steps they prefer assigning to an AI vs. keeping for themselves (F3 from our formative study), we organized these steps into agent and user steps. We then wrote examples of how these steps are composed into plans for tackling particular questions, once again drawing from plans participants wrote in the probe activity. These examples were provided for in-context learning in Cocoa’s system prompt for plan generation (see prompts in Appendix I).
@@
5. Lab Study
We conducted a within-subjects task-based lab study with 16 researchers to evaluate Cocoa and its interactive plan design against a strong chat baseline. Specifically, our study aimed to address the following research questions:
  • L1:
@@ 5. Lab Study
  • L2:
    When did researchers prefer interacting with an AI agent through interactive planning versus our chat baseline?
  • L3:
@@ 5.1. Participants
We recruited 16 Ph.D. and postdoctoral researchers (14 Ph.D.s, 2 postdocs; 10 female, 6 male) in computer science (CS) or CS-adjacent areas via university mailing lists, word of mouth, social media recruitment messages (on Twitter/X, Mastodon, Bluesky), and personal connections. We recruited on a first-come, first-serve basis as we conducted our studies and closed recruitment when we approached data saturation. Further details about our recruitment process, pilot studies, and participants can be found in Appendix F and Table 2 of Appendix E.
@@ 5.2. Baseline System
The baseline system (Figure 11) was a chat interface powered by the same LLM agent used in Cocoa: it used GPT-4o with access to the same tools for completing literature-related tasks using a combination of Semantic Scholar and standard LLM capabilities (summarization, accessing knowledge stored in weights, etc.). While we could integrate interactive plans within chat and vice versa (more on this in Section 8.1), we opted not to for a cleaner comparison of the two design patterns. The design of the chat interface closely resembled that of popular LLM chatbots (e.g., ChatGPT, Claude). During the study, participants positioned the chat interface beside the baseline document editor, which is the same editor that Cocoa uses, except we removed the option to invoke the agent to eliminate all co-planning and co-execution interactions.
@@
Figure 11. Participants’ typical setup for our baseline condition. On the left, the same document editor from Cocoa but without an option to invoke the agent for co-planning and co-execution. On the right, the agent is situated within a chat interface.
[…]
Participants’ typical setup for our baseline condition. On the left, the same document editor from Cocoa but without an option to invoke the agent for co-planning and co-execution. On the right, the agent is situated within a chat interface.
5.3. Research Task and Study Procedure
@@
Because participants identified literature search and understanding as one of the areas they can benefit most from AI assistance (Section 3.2.2), we had participants tackle a literature-augmented task in both Cocoa and the baseline. By literature-augmented, we mean tasks that are to be completed by referencing or drawing from academic literature; in the study, participants worked on not only literature review tasks, but also tasks where researchers needed to make literature-informed decisions such as study design and ideation. This stands in contrast with tasks focused on writing mechanics (e.g., rewording a sentence) or tasks that do not involve literature (e.g., booking travel) that past work has covered (e.g., (Yao et al., 2022a; Yang et al., 2024; Kim et al., 2023; Zhang et al., 2023; Ma et al., 2023; Lee et al., 2024a)).
To make the study more realistic and grounded in real-world research projects, tasks for this study were open questions or unfinished items from participants’ existing project documents. Details for how we controlled for the length of the study and ensured the two tasks were valid and comparable can be found in Appendix G. Tasks were randomly assigned to Cocoa and the baseline.
@@
The first author (study facilitator) conducted 1:1 studies with the 16 participants over Google Meet between October and December 2024. Each study was 90 minutes in length. The study started off with introductions and a brief overview, followed by the two tasks, which were counterbalanced across participants. Detailed procedures for each condition can be found in Appendix [x]. After each condition, participants filled out a short evaluation form with 5-point Likert scale questions for that condition (see Appendix H). The task-based portion of the study took up 75 minutes. The study concluded with a semi-structured interview that lasted around 15 minutes. Participants were asked to reflect on their experiences across the two systems and discuss the pros and cons of both. They were also asked about particular decisions the study facilitator observed when using each system.
Participants received a \$75 USD honorarium upon completion of the study. All studies were recorded and transcribed by Google Meet. This study was reviewed and approved by our organization’s internal IRB.